• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 15:10:02 on February 6,2011

    Entry number 346092

    keyword=DAQ is working, use test17 and Hand

    work done by R. Michaels, L. Selvy, V. Sulkosky, D. Anez

    The DAQ is working now, with some qualifications.

    Simple instructions:

    startcoda Bbite .... and then select test17 (usually it will
    be the choice since it was what ran last)

    startcoda Hand .... and then select HANDalone.

    (Bbite means BBite + L-HRS + R-HRS but no Hand).
    (Hand means HAND alone)


    1. The SFI Aux card on Hand corrupts the data in such a way to
    cause SYNCH loss in EB. I don't want to explain the many hours
    of isolating the problem to this card. Yes, it's the card.
    We tried to find a spare card, but could not. I will hunt for it
    Monday morning, perhaps Cuevas has one. After that we can put
    Hand into the main DAQ.

    2. Various things were wrong with the handshaking on L-HRS and
    R-HRS, not worth documenting here. It works now. The next step is
    to run with HRS singles and do the retiming. Then start checking
    cosmics data on the two HRS. I suppose one could check cosmics
    on Bigbite simultaneously ....

    3. A crew of people are working on Hand as I write this ...